OpenStudy (sphott51):

Can someone help me with graphing absolute value? help with the general form of absolute value of what they are too? y=a|x-b|+c??

OpenStudy (sphott51):

the domain and range? the x and y chart? parent graph? I wasn't in class when they filled it out, I had to take the quiz !

OpenStudy (danjs):

first isolate the absolute value part by itself \[\left| x - b \right| = \frac{ c }{ a }\]

OpenStudy (danjs):

remember the absolute value will give you the positive distance, so (x-b) can be either positive or negative c / a

OpenStudy (danjs):

solve those 2 equations

OpenStudy (danjs):

x - b = c / a or x - b = -c / a